Copper Mountain Group​ (U.S.). The Copper Mountain​ Group, a private equity firm headquartered in​ Boulder, Colorado, borrows £5,200,000 for one year at 8.125​%

interest.

a. What is the dollar cost of this debt if the pound depreciates from $2.0240​/£ to $1.9460​/£ over the​ year?

b. What is the dollar cost of this debt if the pound appreciates from $2.0240​/£ to $2.1680​/£ over the​ year?
Answer #1

Amount to be repaid in Pound = 5,200,000(1+8.125%) = Pound 5,622,500

Dollar received when loan taken = 5,200,000*2.0240 = $10,524,800

a.Dollar to be paid = 5,622,500*1.9460 = $10,941,385

Dollar cost = $416,585 i.e. 3.96%

b.Dollar to be repaid = 5,622,500*2.1680 = $12,189,580

Dollar cost = $1,664,780 i.e. 15.82%